OpenStudy (anonymous):

(2ax+x^2)dy/dx = a^2+2ax solve this differential equations

sam (.sam.):

Variable separable \[\frac{dy}{dx}=\frac{a^2+2 a x}{x (2 a+x)}\] \[y\text{ = }\int\limits \frac{a^2+2 a x}{x (2 a+x)} \, dx\]

sam (.sam.):

Using partial fractions you will get \[\int\limits \left(\frac{a}{2 x}+\frac{3 a}{2 (2 a+x)}\right) \, dx\] Then integrate as usual.

OpenStudy (anonymous):

in second step i doesn't understand

sam (.sam.):

You mean you don't know how to get \[y\text{ = }\int\limits\limits \frac{a^2+2 a x}{x (2 a+x)} \, dx ~~~~~?\]

OpenStudy (anonymous):

ya

sam (.sam.):

From here \[(2ax+x^2)\frac{dy}{dx} = a^2+2ax\] Divide both sides by \(2ax+x^2\) \[\frac{dy}{dx} = \frac{a^2+2ax}{(2ax+x^2)}\] Multiply dx \[dy = \frac{a^2+2ax}{(2ax+x^2)}dx\] \[y = \int \frac{a^2+2ax}{(2ax+x^2)}dx\]

OpenStudy (anonymous):

can you tell me about partial fractions